A slider in a machine moves along a fixed straight rod. Its distance x(in cm) along the rod is given at various times t (in secs).

T : 0 0.1 0.2 0.3 0.4 0.5 0.6

x : 30.28 31.43 32.98 33.54 33.97. 33.48 32.13

Evaluate dx/dt at t = 0.1 And. at t = 0.5


1
Expert's answer
2021-01-22T13:48:59-0500
"\frac {dx}{dt}|_{t=0.1}=\frac {\varDelta x}{\varDelta t}=\frac {31.43-30.28}{0.1}=11.5"


"\frac {dx}{dt}|_{t=0.5}=\frac {\varDelta x}{\varDelta t}=\frac {33.48-33.97}{0.1}=-4.9"